unit digit of a sum
11^11 + 14^14 + 16^16 whats the unit digit of this sum

Answer by drk(1908)

Look for patterns:
11^1 = 11
11^2 = 121
11^3 = 1331
I see the pattern of ending in 1.
---
14^1 = 14
14^2 = 196
14^3 = 2744
when the power is odd it ends in 4, and when it is even it ends in 6.
---
16^1 = 16
16^2 = 256
16^3 = 4096
The units digit is always 8.
---
we get 1 + 6 + 6 = 13
The final units digit will be 3.